# NaVIDA: Vision-Language Navigation with Inverse Dynamics Augmentation

### v1

@@ 7.2 Training Hyperparameters
7.3 Evaluation Metrics Details
@@
7.4 Real-World Deployment Details
Fig. 8 shows the overall architecture of our real-world deployment system. Specifically, NaVIDA was deployed on a remote computation server equipped with an NVIDIA RTX 4090 GPU. The system follows a client–server structure, where the robot functions as the client that collects sensory data and executes movement commands, while the remote server is responsible for high-level reasoning and action prediction.
During operation, the robot continuously captures RGB images from its front-facing camera (as shown in  Fig. 9) and receives natural language or voice commands from the user through a local area network. The sensory data and corresponding commands are then packaged and transmitted to the remote server via the internet. On the server side, the NaVIDA model processes these inputs, performs vision-language reasoning, and predicts the next navigation action. The predicted motion command is then sent back to the robot, where it is converted into executable parameters such as linear velocity, angular velocity, and movement duration through the robot’s motion control API.
@@ 7.4 Real-World Deployment Details
Figure 8: Architecture of the deployed system used for real-world evaluation. RGB frames and user instructions are transmitted from the robot to a remote server running NaVIDA, which predicts action sequences and returns motion commands for execution through the robot’s motion API.
Figure 9: Robot platforms and sensing configuration. AgiBot X2 (left) equipped with a single rectified fisheye RGB camera and Unitree G1 (right) with an additional Intel RealSense D455f front-facing camera, providing consistent monocular RGB input for NaVIDA’s navigation policy.
(a) Distribution of action entropy when truncated@chunk 1
(b) Distribution of action entropy when truncated@chunk 2
(c) Distribution of action entropy when truncated@chunk 3
Figure 10: Action-type entropy across decoding steps under fixed truncation@chunk 1,2,3. Entropy is computed over action-type tokens only. In theory it should increase with horizon; a local drop signals unreliable long-range predictions, motivating truncation before the drop and yielding a more stable executed prefix.
@@
8 More Experiment Results
8.1 More Ablations
@@
Data Scale.  Fig. 11 illustrates the performance trend of NaVIDA on R2R-CE Val-Unseen as a function of the total data size, which includes both IDS and VLN data. As the total data size increases, NaVIDA consistently achieves better performance. The improvement is particularly pronounced when the data scale is small, indicating that NaVIDA effectively leverages additional data to learn more robust representations in low-data regimes. However, as the total data scale continues to grow, the performance gains gradually saturate. This behavior likely reflects a diminishing returns effect, where the model has already captured most of the useful patterns from the available data, and further increases in data size contribute less to performance. Additionally, the stabilization may indicate that other factors such as model capacity, task complexity, or inherent noise in the data become the limiting factors once a certain data threshold is reached.
@@
Figure 11: Navigation performance of NaVIDA on VLN-CE R2R Val-Unseen split with the increase of data scale. SR/OSR/SPL increase and NE decreases as data grows, with gains tapering at larger scales.
@@
Vision Encoder: Frozen vs. Trainable. We conduct an ablation study to evaluate the impact of fine-tuning the vision encoder on overall navigation performance. As shown in  Tab. 8, unfreezing the vision encoder and training it end-to-end provides only marginal improvements compared to keeping it frozen. This result suggests that the pre-trained visual representations are already sufficiently informative for downstream navigation tasks, and further adaptation offers limited benefit. Moreover, fine-tuning substantially increases GPU memory usage and training time, with minimal returns in accuracy. Therefore, we keep the vision encoder frozen in all subsequent experiments, enabling the model to efficiently utilize robust pre-trained visual features while maintaining a favorable balance between performance and computational efficiency.
@@
Number of History Frames. We perform an ablation study to investigate how the number of historical frames influences navigation performance on the R2R-CE Val-Unseen split. All models are trained solely on the R2R training data for consistency. As shown in Tab. 9, using $8$ history frames is sufficient to capture most instruction horizons, providing the model with adequate temporal context for action reasoning. Increasing the number of frames to $16$ does not yield noticeable performance gains and occasionally leads to a slight degradation, likely due to redundant or noisy visual information accumulated over longer sequences. This behavior aligns with findings from NaVILA [8], indicating that excessive historical context can dilute the relevance of recent observations and hinder temporal grounding. Hence, we adopt $8$ history frames as a balanced choice between contextual richness and model efficiency.
@@
Inverse vs. Forward Dynamics. Considering the slower convergence of Forward Dynamics Supervision (FDS), we conducted extended training to further compare its effectiveness and efficiency against Inverse Dynamics Supervision (IDS). Specifically, FDS was trained for a total of $3$ epochs, with checkpoints saved after each epoch for evaluation. As shown in Tab. 10, the performance of FDS gradually improves with additional training epochs, eventually surpassing the baseline model without any dynamic supervision at the third epoch. However, even after prolonged training, FDS still underperforms compared to IDS, which achieves superior results after only a single epoch. This observation highlights that IDS provides more stable and data-efficient supervision signals by directly modeling the causal relationship between consecutive observations and actions. In contrast, FDS requires longer optimization to learn reliable predictive mappings for future states. Overall, these findings underscore the superior training efficiency and effectiveness of IDS in enhancing visual-language navigation performance.
[…]
Base 5.29 61.9 52.6 47.1
    w/ FDS (epoch-1) 6.13 56.0 44.6 38.7
    w/ FDS (epoch-2) 5.06 60.1 51.0 45.2
    w/ FDS (epoch-3) 5.06 62.4 53.9 48.2
    w/ IDS 4.91 63.2 55.5 49.6
Action Entropy  Fig. 10 shows action-entropy distributions under different fixed truncation policies for the execution horizon. We compute action entropy at step $t$, which serves as a proxy for behavioral uncertainty: high $H_{t}$ indicates the policy is uncertain about which action type to choose, while low $H_{t}$ indicates confident selection.
In sequential prediction, a gradual increase of $H_{t}$ with horizon is expected, because uncertainty compounds as visual evidence becomes less informative and small pose errors accumulate. Occasionally, however, we observe local decreases of entropy in later chunks during inference. Such entropy dips often reflect spurious overconfidence induced by dataset bias rather than genuine visual or spatial reasoning. For example, after two consecutive “move forward 75 cm” actions, the training data may frequently contain a “turn right 30 degree” action, which encourages the model to snap to that action with high confidence based on co-occurrence patterns in the text space, even when current images do not support this choice. This kind of shortcut leads to low entropy that is unjustified by the scene and degrades reliability.
These observations motivate an entropy-guided execution-horizon strategy. We retain the prefix that follows the expected rising-entropy trend and truncate before the first entropy dip, that is, at the preceding local peak. Executing only up to this peak avoids high-risk segments dominated by biased priors, reduces error compounding, and improves trajectory stability. Empirically, when truncation is fixed after chunk 1 or 2, the entropy peak aligns with the chosen boundary (Figs. 10(a) and 10(b)); when truncation is disabled and all three chunks are executed, the entropy profile becomes irregular because the unreliable final chunk perturbs the distribution (Fig. 10(c)). This rule detects the earliest deviation from a monotonic rise and stops at the last reliable peak, balancing action completeness and reliability without additional training.
@@
8.2 More Qualitative Results
We present extensive qualitative results of NaVIDA in both real-world and simulated environments to further demonstrate its effectiveness and generalization ability, as shown in Fig. 12Fig. 13, and Fig. 14. The results show that NaVIDA produces smooth, instruction-aligned trajectories in simulation and robustly executes navigation commands in real-world settings. These findings highlight the model’s strong cross-domain generalization, accurate language grounding, and reliable decision-making in complex, long-horizon navigation tasks.

### v2

@@ 0.B.2 Training Hyperparameters
0.B.3 Evaluation Metrics Details
@@
0.B.4 Real-World Deployment Details
Fig.˜7(a) shows the overall architecture of our real-world deployment system. Specifically, NaVIDA was deployed on a remote computation server equipped with an NVIDIA RTX 4090 GPU. The system follows a client-server structure, where the robot functions as the client that collects sensory data and executes movement commands, while the remote server is responsible for high-level reasoning and action prediction.
During operation, the robot continuously captures RGB images from its front-facing camera (as shown in  Fig.˜7(b)) and receives natural language or voice commands from the user through a local area network. The sensory data and corresponding commands are then packaged and transmitted to the remote server via the internet. On the server side, the NaVIDA model processes these inputs, performs vision-language reasoning, and predicts the next navigation action. The predicted motion command is then sent back to the robot, where it is converted into executable parameters such as linear velocity, angular velocity, and movement duration through the robot’s motion control API.
@@ 0.B.4 Real-World Deployment Details
(a) Architecture of the deployed system used for real-world evaluation. RGB frames and user instructions are transmitted from the robot to a remote server running NaVIDA, which predicts action sequences and returns motion commands for execution through the robot’s motion API.
(b) Robot platforms and sensing configuration. AgiBot X2 (left) equipped with a single rectified fisheye RGB camera and Unitree G1 (right) with an additional Intel RealSense D455f front-facing camera, providing consistent monocular RGB input for NaVIDA’s navigation policy.
Figure 7: Real-world evaluation setup.
@@
Appendix 0.C More Experiment Results
0.C.1 More Ablations
@@
Data Scale.  Fig.˜8 illustrates the performance trend of NaVIDA on R2R-CE Val-Unseen as a function of the total data size, which includes both IDS and VLN data. We also constructed a validation set for an IDS on R2R-CE Val-Unseen and plotted the variation curve of the IDS score. As the total data size increases, NaVIDA consistently achieves better performance. The improvement is particularly pronounced when the data scale is small, indicating that NaVIDA effectively leverages additional data to learn more robust representations in low-data regimes. However, as the total data scale continues to grow, the performance gains gradually saturate. This behavior likely reflects a diminishing returns effect, where the model has already captured most of the useful patterns from the available data, and further increases in data size contribute less to performance. Additionally, the stabilization may indicate that other factors such as model capacity, task complexity, or inherent noise in the data become the limiting factors once a certain data threshold is reached.
@@
Figure 8: Navigation performance of NaVIDA on VLN-CE R2R Val-Unseen split with the increase of data scale. SR/OSR/SPL/IDS Score increase and NE decreases as data grows, with gains tapering at larger scales.
@@
Vision Encoder: Frozen vs. Trainable. We conduct an ablation study to evaluate the impact of fine-tuning the vision encoder on overall navigation performance. As shown in  Tab.˜8, unfreezing the vision encoder and training it end-to-end provides only marginal improvements compared to keeping it frozen. This result suggests that the pre-trained visual representations are already sufficiently informative for downstream navigation tasks, and further adaptation offers limited benefit. Moreover, fine-tuning substantially increases GPU memory usage and training time, with minimal returns in accuracy. Therefore, we keep the vision encoder frozen in all subsequent experiments, enabling the model to efficiently utilize robust pre-trained visual features while maintaining a favorable balance between performance and computational efficiency.
@@
Number of History Frames. We perform an ablation study to investigate how the number of historical frames influences navigation performance on the R2R-CE Val-Unseen split. All models are trained solely on the R2R training data for consistency. As shown in Tab.˜9, using $8$ history frames is sufficient to capture most instruction horizons, providing the model with adequate temporal context for action reasoning. Increasing the number of frames to $16$ does not yield noticeable performance gains and occasionally leads to a slight degradation, likely due to redundant or noisy visual information accumulated over longer sequences. This behavior aligns with findings from NaVILA [navila], indicating that excessive historical context can dilute the relevance of recent observations and hinder temporal grounding. Hence, we adopt $8$ history frames as a balanced choice between contextual richness and model efficiency.
@@
0.C.2 More Qualitative Results
We present extensive qualitative results of NaVIDA in both real-world and simulated environments to further demonstrate its effectiveness and generalization ability, as shown in Fig.˜9Fig.˜10, and Fig.˜11. The results show that NaVIDA produces smooth, instruction-aligned trajectories in simulation and robustly executes navigation commands in real-world settings. These findings highlight the model’s strong cross-domain generalization, accurate language grounding, and reliable decision-making in complex, long-horizon navigation tasks.